Consider an economy in long run equilibrium described by the following equations: Y = C + I + G + NX Y = 5000 G = 1000 T = 1000 C = 250 + 0.75*( Y - T ) I = 1000 - 50*r NCO = 500 - 50*r Where r is the real interest rate (in % terms). Suppose G rises to 1250 without any change in T. Solve again for the equilibrium real interest rate and the rest of the variables you found in (b). Explain your results. i.) The new equilibrium real interest rate in this economy is _______%. A. 2.5 B. 5 C. 10 D. 7.5 The new equilibrium national savings for this economy is __________. A. 600 B. 750 C. 850 D. 500 The new equilibrium investment for this economy is _________ , while new equilibrium net capital outflow (NCO) is ________. A. 600 ; -100 B. 400 ; 100 C. 500 ; 100 D. 500 ; 0 The new equilibrium level of the trade balance (NX) in this economy is ______. A. 150 B. 100 C. 250 D. 0 Relative to the closed economy case, this increase in the budget deficit causes ______ crowding out of domestic investment because the ________ in the real interest rate causes financial capital to __________ the country. A. More ; increase ; flow out of B. More ; decrease ; flow into C. Less ; increase ; flow out of D. Less ; increase ; flow into
Answer #1

S= 5,000-3,250-1,000 = 750 I 1,000-250 750 1000-50*5 S-I=NX=0 (750-750) 0 = 500-500e,-> ε = 1 a) c-250+0.75(5000-1000):3250